OpenStudy (anonymous):

Find a quadratic model for the set of values: (-2, -20), (0, -4), (4, -20). Show your work. please help!

OpenStudy (amistre64):

well, we can go a few different ways on this

OpenStudy (anonymous):

the easiest please!

OpenStudy (amistre64):

oh there is no 'easiest' there is logical, there is systematic, there is matrix .... there is no easiest

OpenStudy (anonymous):

please just explain a way to solve this.

OpenStudy (amistre64):

logical set up coefficients such that for every x value, the rest of them zero out y = a + b(x+2) +cx(x+2) now when x=-2, y=-20 -20 = a + b(-2+2) +c(-2)(-2+2) b and c goto 0 leaving us -20 = a ------------------------ next, let x=0, y=-4 -4 = a + b(0+2) +c(0)(0+2) c goto 0 leaving us -4 = -20 +2b, b=16/2 = 8 ------------------------- now let x = 4, y=-20; and c is the only thing left to determine -20 = -20 + 8(4+2) +c(4)(4+2)

OpenStudy (amistre64):

systematicly we know ax^2 +bx +c = y gives us a quadratic; setup a system of 3 equation and the 3 unknowns can be determined (-2)^2 a + (-2) b + c = -20 ( 0)^2 a + ( 0) b + c = - 4 ( 4)^2 a + ( 4) b + c = -20 c = -4 by default and this reduces to a 2 system setup 4a -2b = -20-4 16a +4b = -20-4

OpenStudy (amistre64):

-20+4 that is

OpenStudy (amistre64):

matrix is just to take the coeffs of the system .. and row reduce it to echelon form (rref)